Exercise 8-11A Determining materials price and usage variances LO 8-6

Monique’s Florals produced a special Mother’s Day arrangement that included eight roses. The standard and actual costs of the roses used in each arrangement follow:

Standard Actual
Average number of roses per arrangement 7.10 7.50
Price per rose × $ 0.50 × $ 0.44
Cost of roses per arrangement $ 3.55 $ 3.30

Monique’s Florals planned to make 820 arrangements but actually made 930.

Required

  1. Determine the total flexible budget materials variance and indicate whether it is favorable (F) or unfavorable (U).

  2. Determine the materials price variance and indicate whether it is favorable (F) or unfavorable (U).

  3. Determine the materials usage variance and indicate whether it is favorable (F) or unfavorable (U).

  4. Confirm the accuracy of Requirements a, b, and c by showing that the sum of the price and usage variances equals the total variance.
Answer #1

a) Total flexible budget material variance = Standard cost-Actual Cost

                                                        = (3.55-3.30)*930

Total flexible budget material variance = 232.50 F

b) Material price variance = (0.50-0.44)*6975 = 418.5 F

c) Material quantity variance = (7.1-7.5)*930*.50 = 186 U

d) Sum of price and usage variance = 418.50 F+186 U = 232.50 F